OpenStudy (anonymous):

sin(tan^-1x-sin^-1x)express algebraically in terms of x please need help with this and please explain each step

OpenStudy (neer2890):

you mean \[\sin (\tan^{-1} x-\sin^{-1} x)\]???

OpenStudy (anonymous):

yes that is the sum

OpenStudy (neer2890):

\[\sin (\frac{ \sin^{-1} x }{ \cos^{-1} x }- \sin^{-1} x)\] \[\sin (\frac{ \sin^{-1} x(1- \cos^{-1} x)}{ \cos^{-1} x })\] \[x(\frac{ 1 }{ \cos^{-1} x }-1)\] \[x(\sec ^{-1}x-1)\]

OpenStudy (anonymous):

can you just explain the third step please

OpenStudy (neer2890):

\[\sin*\sin ^{-1}x(\frac{ 1-\cos^{-1} x }{ \cos^{-1} x })\] now sin and sin inverse cancels out and x remains